OpenStudy (therealmeeeee):

Please help :P Solve x + 7 = 13. A. –20 B. 20 C. –6 D. 6

OpenStudy (anonymous):

Well, 13-7=6. So the answer is D. 6. May I please have a fan and a medal? ;)

OpenStudy (therealmeeeee):

I have more can you help me??

OpenStudy (anonymous):

Sure!

OpenStudy (therealmeeeee):

Solve 11 – x = 3. A. 14 B. –8 C. 8 D. –14

OpenStudy (therealmeeeee):

I will make another Post so you can get more medals

OpenStudy (anonymous):

Ok but the answer to this question is C. 8 because 11-3=8. ;)

OpenStudy (therealmeeeee):

Thanks again xD

OpenStudy (anonymous):

Do you need help with any more questions?